A city has an equilibrium price for new housing of $100,000. The price elasticity of demand...

  1. A city has an equilibrium price for new housing of $100,000. The price elasticity of demand for new housing is 1 and the price elasticity of supply is 4. The city imposes a development tax of $25,000 per house. Show the effects of the development tax on the housing markets:
    1. What are the housing prices in the city with and without the tax?
    2. What is the marginal cost of production with and without the tax?
    3. What is the relative incidence of the tax (with respect to demand and supply)?
Answer #1

More elastic segment of the market bears larger tax burden, so consumers bear more burden

Thus consumers bears 4 times the burden , borne by producers .

So 1) without tax, housing price = 100,000

with tax, = 120,000

2) MC of production without tax = 100,000

with tax = 105,000

3) total tax burden on buyers = 20,000/25,000 = 4/5 =.8

burden on sellers = 5,000/25,000 = 1/5 = .2